OpenStudy (anonymous):

what is the y-intercept of the line that goes through the points (6,14) and (18,8)?

OpenStudy (anonymous):

over twelve down six slope is \[-\frac{1}{2}\]

OpenStudy (anonymous):

now use almighty point slope formula \[y-y_1=m(x-x_1)\] with \[x_1=6,y_1=14, m=-\frac{1}{2}\]

myininaya (myininaya):

first lets write the equation first y=mx+b we need m and b \[m=\frac{ \Delta y}{\Delta x}=\frac{8-14}{18-6}=\frac{-6}{12}=\frac{-1}{2}\] y=-1/2*x+b we know a point on this line (x=6,y=14) 14=-1/2*6+b 14=-3+b 17=b y=-1/2*x+17 y intecept is b=17 x intercept (let y=0 solve for x) 0=-1/2*x+17 17=1/2*x 2*17=x 34=x

OpenStudy (anonymous):

typing as fast as i can. hello my myininaya!

OpenStudy (anonymous):

\[\color{red}{\text{i tried to tease you the other day in a post congratulating amistre.}}\]